Part 2: For the following problems solve the IVP - be careful of initial conditions and coefficients which change in each pro1.8 y + 5y + 6y = 5e-5t y(0) = 0, y(0) = 0 1.9 y + 6y = t y(0) = 1

Please solve using LAPLACE transforms and inverses. Will upvote if correct!!!
Answer #1

ylc)=-4, 9021 6 1.5) yh+6y +84 -0, L {yt byl +89] = {0} Lly]+62{y}+8 L {y}=0 SAL14] -590-10) +6 {SL {41 - 40)] +81142-0 Sན S S S - 117) y + 6y +8y=5 y0)=, Y00)=1 Llyll+byl+8 y} = 2158 s? L{y}-5.40)-410) +6[5L{9}–400)}+84493 = „ s²LhY}-s-it GS